# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v2

@@ 1 Introduction
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. However, existing researches often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks. Additionally, most of these studies tend to assume a fixed route travel time, excluding potential queuing delays on the road. As a result, their outcomes may become distorted, as the actual travel time of the robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction. Second, the development of a precise model to quantify system efficiency within the RSS is essential. Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity. Third, we emphasize the need for management strategies that leverage the flexibility of RSS to achieve cost-effective responses to fluctuating demands. These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
@@ 4.1 Virtual platoon and cycle
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network perimeter control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed information on the RC framework, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput. CA* is based on a simple prioritized-planning scheme: each agent is assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
@@
We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets. The RC-S parameters, detailed in Table 2, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 10 rounds of replication for each test. In each experiment, we set a warm-up period of 10 minutes and then record the system performance for 50 minutes of continuous operations.
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail. We conduct 20 rounds of replication for each scenario. In each experiment, we set a warm-up period of one hour and then record the system performance for three hours of continuous operation.
@@ 6.3 Efficiency underestimating in queueing models
Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (Buzacott and Shanthikumar,, 1993). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing. However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3. Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2. The results are shown in Table 5. The CQN model significantly underestimates system throughput in scenarios with few robots. It fails to capture variations in traffic congestion across different flow rates and overestimates road conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
Table 5: Error in throughput estimation
@@
RC-S 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
@@ 7 Layout Design Optimization
8 Numerical Examples
@@
8.1 Site rental cost
@@
8.2 Labor cost

### v3

@@ 1 Introduction
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. Zou et al., (2021) proposed several closed queuing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput. Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express. However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts. This aspect makes them less suited for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper. On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks. Consequently, their results may become distorted as the actual travel time of robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions. Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to clarify the relationship between system configurations and sorting capacity. Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands. These contributions aim to improve the performance and scalability of RSS implementations.
@@ 4.1 Virtual platoon and cycle
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed information on the RC framework, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput. CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
@@
We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets. The RC-S parameters, detailed in Table 2, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 50 repeated experiments for each scenario. In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operations. Our algorithms and simulations are coded in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail.
@@ 6.3 Efficiency underestimating in queueing models
Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (John A. Buzacott,, 1993). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing. However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3. Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2. The results are shown in Table 5, where Eq.(28) represents the proposed estimation formula. The CQN model significantly underestimates system throughput in scenarios with few robots. It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
Table 5: Error in throughput estimation under RC-S control
@@
Eq.(28) 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
we also conduct a experiments on a real case of Deppon Express described in Zou et al., (2021), where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations and 170 robots. The The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ were set according to the data provided by Deppon Express. We compared the performance of RC-S with the results from Zou et al., (2021), as shown in Table 6, where $T_{D}$, $T_{CQN}$ and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under RC-S scheme, respectively. Combined with the experimental results in section 6.1, RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput. Meanwhile, the CQN model effectively captures the congestion in the RSS system implemented by Deppon Express, but underestimates the traffic capacity of the network when the RC-S scheme is applied.
Table 6: Real Case Validation
$D(m)$ $r_{w}(/s)$ $v_{max}(m/s)$ $\tau_{c}(s)$ $v_{VP}(m/s)$ $T_{D}(/h)$ $T_{CQN}(/h)$ $T_{RCS}(/h)$
0.6 0.42 2 1.2 2 7,163 6,907 8,758
6.4 Evaluating RC-S across diverse scenarios
In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not spread across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits. The placement of outlets is also restricted by site factors, such as load-bearing columns or the presence of certain equipment. Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across each outlet. In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation. For each set of experiments, we evaluate the overall performance across 50 test runs. Each run includes a 30-minute warm-up period, followed by an observation period where average data is recorded over one hour.
We first compared the impact of different numbers of active loading stations on average throughput. We consider five scenarios with varying network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for sorting process. Other RC-S parameters are specified in detail in Table 2. Figure 16 shows the trend of throughput across the five scenarios. It can be observed that large-scale networks not only have a higher throughput upper bound, but their slope—representing the average efficiency per loading station—is also generally higher. This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
Figure 16: The effect of different numbers of active loading stations
(a) Division of outlet areas in the network
(b) RC-S performance in different scenarios
Figure 17: The effect of closing outlets in the network
We next evaluate the effect of closing some outlets on system throughput. The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters remaining constant. We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center. To control variates, the number of closed outlets in each scenario is set to 1/9 of the total. Scenarios (2), (3), and (4) correspond to areas 1, 2, and 3 in Figure 17(a), respectively. Results from 50 repeated experiments for each scenario are shown in Figure 17(b). The green dashed line represents the average throughput when all outlets are open. It can be observed that randomly closing outlets has no significant impact on system throughput. Closing outlets in areas 1 and 2 results in a decrease in average throughput, while closing outlets in area 3 leads to a slight increase in average throughput. It is intuitive that closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network. However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources. Since most shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
Figure 18: The effect of heterogeneous demand distribution
In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency. In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency for each outlet; however, this cannot entirely eliminate demand heterogeneity. We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution. For outlets with extremely low arrival rates, the situation is similar to the previous experiments involving closed outlets, and their impact can be minimized by random assignment. The two types of outlets are spatially uniformly distributed throughout the network. Keeping other experimental parameters constant, we adjust the proportion of the two types of outlets, and the system performance are shown in Figure 18. The results indicate that with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains efficiency, and the impact of demand heterogeneity on system throughput is negligible. This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow across the network.
Table 7: RC-S performance in large-scale system
$n_{h},n_{v}$ Throughput Avg. path Avg. time Max. time # Feasible paths
($\cdot{10}^{4}$/h) length (m) per cycle (s) per cycle (s)
22 3.871 27.832 0.0359 0.0942 118,136
24 4.331 30.062 0.0480 0.1255 167,084
26 4.699 32.517 0.0614 0.2393 229,888
28 5.156 34.734 0.0836 0.2436 308,948
30 5.517 37.261 0.1004 0.2820 406,856
[…]
@@ 7 Layout Design Optimization
8 Numerical Examples for Layout Design
@@
8.1 Sensitivity Analysis on the site rental cost
@@
8.2 Sensitivity Analysis on the labor cost